OpenStudy (anonymous):

Choose the possible number of solutions for these equations. Also, can you show me how you chose that solution. 2x - 3y = 12, -x + 3/2y = -6 either: infinite amount of solution no solutions one solution

OpenStudy (anonymous):

infinite

OpenStudy (saifoo.khan):

Right.

OpenStudy (saifoo.khan):

coz the lines are on each other.

OpenStudy (anonymous):

second equation: -x+(3/2)y=-6 -2x+3y=-12 2x-3y=12 thus second equation is same as first equation therefore infinite solutions